# Causal2Vec: Improving Decoder-only LLMs as Versatile Embedding Models

### v1

@@ 3.1 The Contextual Token
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the vertical concatenation operation, $\boldsymbol{\mathrm{I}}\in\mathbb{R}^{l\times d}$ and $\boldsymbol{\mathrm{T}}\in\mathbb{R}^{n\times d}$ represent the word embeddings of the task-specific instruction and input text, respectively. In this way, each token following the Contextual token $C$ can capture contextualized information even without attending to future tokens. More importantly, the use of Contextual token requires no modifications to the model architecture, which not only preserves LLMs’ ability to extract semantic information learned during pretraining, but also enables seamless integration across different LLMs.
@@ 4.4 MTEB Results
Comparison to Bidirectional LLM-based Methods. LLM2Vec (llm2vec, ) removes the causal attention mask to enable bidirectional attention, demonstrating competitive performance on the MTEB benchmark. NV-Embed (nvembed, ) builds upon this approach by introducing a latent attention layer to obtain pooled embeddings, addressing the limitations of mean pooling. Notably, NV-Embed achieves significant improvements by incorporating additional non-retrieval MTEB-related and synthetic data. However, when trained solely on public retrieval data, NV-Embed shows lower performance compared to LLM2Vec. GRITLM (grit, ) unifies embedding and generation training paradigms, yielding embedding performance on par with LLM2Vec. In contrast to these bidirectional LLM-based methods, our Causal2Vec requires no modifications to the model architecture, yet enables each token in the sequence to access contextual information through the introduced Contextual token. More importantly, our method consistently outperforms LLM2Vec across various base models under identical training data, achieving improvements of 0.780.780.78 points for S-LLaMA-1.3B, 0.800.800.80 for LLaMA-2-7B, and 1.301.301.30 for Mistral-7B. These results underscore that shifting from causal attention to bidirectional attention is not necessary for adopting LLMs to text embedding tasks—and may even compromise the model’s ability to extract the well-learned semantic information. We argue that the effectiveness of bidirectional attention in capturing contextual information relies on maintaining consistency in attention mechanism throughout both pretraining and finetuning.
@@
Comparison to ECHO. ECHO echo repeats the input, allowing each token from the second occurrence to access the complete sequence content. However, this strategy comes with a significant drawback: it doubles the maximum sequence length, thus increasing computational cost during both training and inference. Conversely, our Causal2Vec enables contextual information access through a single Contextual token, without introducing additional overhead to the LLM. Furthermore, our proposed representation method, which concatenates the the last hidden states of Contextual and EOS tokens as final text embedding, effectively mitigates the recency bias inherent in last-token pooling used by ECHO and reduces reliance on the EOS token. As a result, Causal2Vec consistently outperforms ECHO on S-LLaMA-1.3B and Mistral-7B by 0.620.620.62 points and 1.421.421.42 points, confirming the effectiveness of our design in leveraging LLMs’ causal attention for text embedding.
@@
Comparison to bge-en-icl. bge-en-icl (bgeicl, ) endows LLM-based embedding models with in-context learning (ICL) (icl, ) capabilities by incorporating multiple task-related examples into the input. Our method significantly outperforms bge-en-icl (few-shot) on clustering, reranking, and STS tasks with an improvement of at least 1.691.691.69 points, while yielding comparable results on pair classification and summarization tasks. This indicates that our method achieves stronger overall performance on unseen non-retrieval tasks and generates universal text embeddings without relying on task-related demonstrations. More importantly, bge-en-icl relies on knowledge distillation from a reranker model liu2025matryoshka as the teacher—an effective strategy that is not adopted by other baselines. In addition, incorporating in-context examples substantially increases the computational burden of LLMs, especially given that the maximum sequence length of bge-en-icl can reach up to 2048 tokens, which is four times that of our method, significantly limiting its applicability in resource-constrained scenarios. Under the compute-matched setting, our model achieves an average score improvement of 1.431.431.43 over bge-en-icl (zero-shot). Moreover, bge-en-icl is sensitive to the selection strategy and number of in-context examples, hindering its applicability to different LLMs, especially for language models that lack in-context learning capabilities. Notably, our method is orthogonal to bge-en-icl. When equipped with the same ICL strategy, Causal2Vec (w/ ICL) achieves a state-of-the-art average score of 66.85, outperforming bge-en-icl (few-shot) by 0.77 points.
@@ 4.5.1 Effectiveness of Each Component
To evaluate the effectiveness of the proposed Contextual token and representation method, we conduct ablation studies on MTEB-MINI using two base models of different scales: S-LLaMA-1.3B and Mistral-7B. As shown in Table 2, incorporating the Contextual token into LLM’s causal attention mechanism yields average score improvements of 0.240.240.24 and 0.370.370.37 for S-LLaMA-1.3B and Mistral-7B, respectively. These results not only confirm the effectiveness of the Contextual token but also highlight its scalability and applicability across different LLMs. We attribute these performance improvements to the rich contextualized content encoded in the Contextual token, which allows preceding tokens in the sequence to access accurate sentence information even without attending to future tokens, thereby mitigating the inherent architectural limitation in causal attention.
By concatenating the last hidden states of Contextual and EOS tokens as the final vector representation, we observe consistent performance improvements across all seven tasks, with average score gains of 0.720.720.72 and 0.560.560.56 on S-LLaMA-1.3B and Mistral-7B compared to standard LLMs, respectively. These results further confirm the generalizability and robustness of our method in enhancing embedding quality. We attribute the effectiveness of the proposed representation method to the following reasons: (1) it effectively alleviates the recency bias, as the Contextual token is not influenced by tokens near the end of sequence; (2) it enables explicit supervision of the Contextual token during training; and (3) the concatenation of two context-aware tokens enriches the semantic information of the final text embedding. See Appendix C.2 for further discussion on different representation methods.
@@ 4.5.2 Impact of Freezing the Bidirectional Encoder
Since the BERT-style bidirectional encoder (E5-base-v2) we use is specifically trained for embedding tasks, this section investigates whether it should be frozen during finetuning. As shown in Table 4, we observe that finetuning the bidirectional encoder with LoRA leads to an average score improvement of 0.290.290.29 on the MTEB-MINI for S-LLaMA-1.3B, but degrades Mistral-7B’s performance by 0.130.130.13 points. We attribute this to two potential effects of making the bidirectional encoder trainable: (1) it may cause catastrophic forgetting in the bidirectional encoder, and (2) it may help the LLM better interpret the Contextual token through joint finetuning. Large-scale LLMs are more susceptible to the former, as they already have sufficient capacity to comprehend newly added tokens during finetuning (llava, ; blip, ). This suggests that whether the introduced bidirectional encoder should remain frozen may depend on the scale of the underlying LLM.

### v2

@@ 3.1 The Contextual Token
where $\cdot$$\cdot$] denotes the vertical concatenation operation, $\boldsymbol{\mathrm{I}}\in\mathbb{R}^{l\times d}$ and $\boldsymbol{\mathrm{T}}\in\mathbb{R}^{n\times d}$ represent the word embeddings of the task-specific instruction and input text, respectively. In this way, each token following the Contextual token $C$ can capture contextualized information even without attending to future tokens. More importantly, the use of Contextual token requires no modifications to the model architecture, which not only preserves LLMs’ ability to extract semantic information learned during pretraining, but also enables seamless integration across different LLMs.
@@ 4.4 MTEB Results
Comparison to Bidirectional LLM-based Methods. LLM2Vec (llm2vec, ) removes the causal attention mask to enable bidirectional attention, demonstrating competitive performance on the MTEB benchmark. NV-Embed (nvembed, ) builds upon this approach by introducing a latent attention layer to obtain pooled embeddings, addressing the limitations of mean pooling. Notably, NV-Embed achieves significant improvements by incorporating additional non-retrieval MTEB-related and synthetic data. However, when trained solely on public retrieval data, NV-Embed shows lower performance compared to LLM2Vec. GRITLM (grit, ) unifies embedding and generation training paradigms, yielding embedding performance on par with LLM2Vec. In contrast to these bidirectional LLM-based methods, our Causal2Vec requires no modifications to the model architecture, yet enables each token in the sequence to access contextual information through the introduced Contextual token. More importantly, our method consistently outperforms LLM2Vec across various base models under identical training data, achieving improvements of $0.78$ points for S-LLaMA-1.3B, $0.80$ for LLaMA-2-7B, and $1.30$ for Mistral-7B. These results underscore that shifting from causal attention to bidirectional attention is not necessary for adopting LLMs to text embedding tasks—and may even compromise the model’s ability to extract the well-learned semantic information. We argue that the effectiveness of bidirectional attention in capturing contextual information relies on maintaining consistency in attention mechanism throughout both pretraining and finetuning.
@@
Comparison to ECHO. ECHO echo repeats the input, allowing each token from the second occurrence to access the complete sequence content. However, this strategy comes with a significant drawback: it doubles the maximum sequence length, thus increasing computational cost during both training and inference. Conversely, our Causal2Vec enables contextual information access through a single Contextual token, without introducing additional overhead to the LLM. Furthermore, our proposed representation method, which concatenates the the last hidden states of Contextual and EOS tokens as final text embedding, effectively mitigates the recency bias inherent in last-token pooling used by ECHO and reduces reliance on the EOS token. As a result, Causal2Vec consistently outperforms ECHO on S-LLaMA-1.3B and Mistral-7B by $0.62$ points and $1.42$ points, confirming the effectiveness of our design in leveraging LLMs’ causal attention for text embedding.
@@
Comparison to bge-en-icl. bge-en-icl (bgeicl, ) endows LLM-based embedding models with in-context learning (ICL) (icl, ) capabilities by incorporating multiple task-related examples into the input. Our method significantly outperforms bge-en-icl (few-shot) on clustering, reranking, and STS tasks with an improvement of at least $1.69$ points, while yielding comparable results on pair classification and summarization tasks. This indicates that our method achieves stronger overall performance on unseen non-retrieval tasks and generates universal text embeddings without relying on task-related demonstrations. More importantly, bge-en-icl relies on knowledge distillation from a reranker model liu2025matryoshka as the teacher—an effective strategy that is not adopted by other baselines. In addition, incorporating in-context examples substantially increases the computational burden of LLMs, especially given that the maximum sequence length of bge-en-icl can reach up to 2048 tokens, which is four times that of our method, significantly limiting its applicability in resource-constrained scenarios. Under the compute-matched setting, our model achieves an average score improvement of $1.43$ over bge-en-icl (zero-shot). Moreover, bge-en-icl is sensitive to the selection strategy and number of in-context examples, hindering its applicability to different LLMs, especially for language models that lack in-context learning capabilities. Notably, our method is orthogonal to bge-en-icl. When equipped with the same ICL strategy, Causal2Vec (w/ ICL) achieves a state-of-the-art average score of 66.85, outperforming bge-en-icl (few-shot) by 0.77 points.
@@ 4.5.1 Effectiveness of Each Component
To evaluate the effectiveness of the proposed Contextual token and representation method, we conduct ablation studies on MTEB-MINI using two base models of different scales: S-LLaMA-1.3B and Mistral-7B. As shown in Table 2, incorporating the Contextual token into LLM’s causal attention mechanism yields average score improvements of $0.24$ and $0.37$ for S-LLaMA-1.3B and Mistral-7B, respectively. These results not only confirm the effectiveness of the Contextual token but also highlight its scalability and applicability across different LLMs. We attribute these performance improvements to the rich contextualized content encoded in the Contextual token, which allows preceding tokens in the sequence to access accurate sentence information even without attending to future tokens, thereby mitigating the inherent architectural limitation in causal attention.
By concatenating the last hidden states of Contextual and EOS tokens as the final vector representation, we observe consistent performance improvements across all seven tasks, with average score gains of $0.72$ and $0.56$ on S-LLaMA-1.3B and Mistral-7B compared to standard LLMs, respectively. These results further confirm the generalizability and robustness of our method in enhancing embedding quality. We attribute the effectiveness of the proposed representation method to the following reasons: (1) it effectively alleviates the recency bias, as the Contextual token is not influenced by tokens near the end of sequence; (2) it enables explicit supervision of the Contextual token during training; and (3) the concatenation of two context-aware tokens enriches the semantic information of the final text embedding. See Appendix C.2 for further discussion on different representation methods.
@@ 4.5.2 Impact of Freezing the Bidirectional Encoder
Since the BERT-style bidirectional encoder (E5-base-v2) we use is specifically trained for embedding tasks, this section investigates whether it should be frozen during finetuning. As shown in Table 4, we observe that finetuning the bidirectional encoder with LoRA leads to an average score improvement of $0.29$ on the MTEB-MINI for S-LLaMA-1.3B, but degrades Mistral-7B’s performance by $0.13$ points. We attribute this to two potential effects of making the bidirectional encoder trainable: (1) it may cause catastrophic forgetting in the bidirectional encoder, and (2) it may help the LLM better interpret the Contextual token through joint finetuning. Large-scale LLMs are more susceptible to the former, as they already have sufficient capacity to comprehend newly added tokens during finetuning (llava, ; blip, ). This suggests that whether the introduced bidirectional encoder should remain frozen may depend on the scale of the underlying LLM.